OpenStudy (nsh4267):

A construction firm is trying to decide which of two models of a crane to purchase. Model A costs $98,000 and requires $4,000 per year to maintain. Model B has an initial cost of $60,000 and a maintenance cost of $12,000 per year. For how many years must model A be used before it becomes more economical than B?

OpenStudy (anonymous):

98000 + 4000x = 60000 + 12000x -60k from each side 38000 + 4000x = 12000x -4000x from each side 38000 = 8000x (or 38 = 8x and divide by 8) Divide by 8000 4.75 = x , so 4.75 years